# Contrastive-KAN: A Semi-Supervised Intrusion Detection Framework for Cybersecurity with scarce Labeled Data

### v1

@@ Abstract
In the era of the Fourth Industrial Revolution, cybersecurity and intrusion detection systems are vital for the secure and reliable operation of IoT and IIoT environments. A key challenge in this domain is the scarcity of labeled cyber-attack data, as most industrial systems operate under normal conditions. This data imbalance, combined with the high cost of annotation, hinders the effective training of machine learning models. Moreover, rapid detection of attacks is essential, especially in critical infrastructure, to prevent large-scale disruptions. To address these challenges, we propose a real-time intrusion detection system based on a semi-supervised contrastive learning framework using the Kolmogorov–Arnold Network (KAN). Our method leverages abundant unlabeled data to distinguish between normal and attack behaviors effectively. We validate our approach on three benchmark datasets, UNSW-NB15, BoT-IoT, and Gas Pipeline, using only 2.20%, 1.28%, and 8% of labeled samples, respectively, to simulate real-world conditions. Experimental results show that our method outperforms existing contrastive learning-based approaches. We further compare KAN with a traditional multilayer perceptron (MLP), demonstrating KAN’s superior performance in both detection accuracy and robustness under limited supervision. KAN’s ability to model complex relationships and its learnable activation functions are also explored and visualized, offering interpretability and potential for rule extraction. The method supports multi-class classification and proves effective in safety-critical environments where reliability is paramount.
@@ 1 Introduction
Today with the rapid development of networking, the internet of things (IoT) and the industrial internet of things (IIoT), including technologies such as smart home applications, medical devices, automated vehicles and sensor networks [1], the adoption of networking is accelerating. Cyber-security has become one of the most critical aspects of both cyber-space and networks. The need for secure digital environments is evident in the era of fourth industrial revolution, where internet connectivity is integrated into many aspect of modern life [2], [3], [4].
@@
In IoT and IIoT scenarios, the intrusion detection systems are designed to identify suspicious patterns in sensor, actuator and network packet data. With the advent of machine learning (ML) and deep learning (DL), the field of intrusion detection has shifted from classical rule-based methods towards intelligent data-derive approaches. ML approaches require fewer computational resources and data compared to DL methods. However, these methods often lack generalizability needed to handle unseen data. In contrast, DL methods offer superior performance but require extensive volumes of data and much larger computational resources [2]. To address these challenges, researches have proposed various solutions. Unsupervised methods aim to address the problem of costly and time-consuming data annotation by leveraging unlabeled data [5]. However, these approaches usually lack the desired performance that is needed. For data privacy, federated learning approaches are proposed to train the networks without the need to centralize data [2]. Despite all said, many existing methods still lack the interpretability required by industries for safe and reliable operation [6].
@@
To address data augmentation, labeled data scarcity or data imbalance handling, many algorithms rely on oversampling techniques, such as synthetic minority over-sampling technique (SMOTE) [7], or its variants. Generative appearances using generative adversarial networks (GAN) have also been explored. However, there are a few problems we can point out to rule out the use of such techniques. First, these methods will have performance degradation if the imbalance ratio is large, which is common in intrusion detection datasets. Second, in many of these datasets, the data distributions of attack and normal samples are quite similar. As a result, using oversampling techniques can lead to unclear boundaries between attack and normal cases. And regarding GANs, often the training of these networks is difficult. In this work, the use of such methods for imbalance handling and data augmentation is discouraged. Instead, we use a masking strategy to randomly mask out a certain percentage of the data.
@@
Another commonly used technique in intrusion detection is feature selection [8]. However, one significant problem with feature selection is that if an attack targets a specific feature that has been removed during feature selection, the model may fail to detect it. In this work, we avoid feature selection, but the features with no variations in the Gas Pipeline dataset are removed. We perform no feature selection other than the mentioned one.
@@
One of the major challenges in real-time intrusion detection is the pre-processing that could delay real-time detection of cyber intrusions. Many methods use window-based sampling approaches. the problem with such techniques is that the algorithm must wait for a complete data window, before proceeding further with operations such as averaging or applying one-dimensional convolutional neural networks (1D-CNNs) for feature extraction and final classification. Furthermore the window-based algorithms, often exhibit lower performances [9]. Another preprocessing method used frequently in the literature is principle component analysis (PCA), which reduces the dimensionality of the data or transforms the input data into a set of orthogonal components based on maximum variation along the principle components. Our approach is based on individual data samples, which eliminates the need for windowing. And the pre-processing associated with this method only consists of feature encoding and standard scaling for normalization. This preserves the interpretability of the features and timely intrusion detection.
@@
As discussed before, the limited availability of annotated data is one of the biggest challenges in ML and DL. We mentioned the advantages and disadvantages of ML and DL. To bridge the gap between ML and DL and to address the problem of annotated data, semi-supervised approaches have been proposed where the network is first pre-trained on large amounts of unlabeled data, then using only a small labeled subset, the fine-tuning is performed. One example of these semi-supervised leaning is the SimCLR framework introduced in [10], which we adopt in this work with KAN. In this study, a feature extractor based on a single layer of KAN is trained using the abundant unlabeled data. These data points are first augmented into two different views, and the feature extractor is to create same representations for these augmented views which are literally the different views of the same data point. Once the pre-training is completed, fine-tuning is performed using only a small subset of data with labels.
@@
To address the interpretability and explainability challenges faced by DL models, we incorporate KAN into our framework. First proposed in [11], KAN is a variant of fully connected layer, where the traditional weights are replaced with learnable activation functions in the form of splines. These splines approximate univariate functions that can capture complex patterns in the data. As will be further discussed in later sections, KAN has a better performance compared to the MLP networks with the same number of parameters. KAN decomposes the complex multi-variate inputs into multiple univariate functions, and originally was used for solving ODEs and PDEs [12], but recent studies extend KAN’s applications to time series analysis [13] and classification tasks.
The contributions of the study are listed as follows:
@@
  • An efficient real-time intrusion detection system based on semi-supervised contrastive learning, utilizing large amounts of unlabeled data.
@@
  • Utilizing a small fraction of labeled data and achieving performance better than other state-of-the-art methods.
@@
  • A thorough ablation study and comparison to traditional MLP networks that explores different hyper-parameters of the KAN model to determine which hyper-parameters have the most impact on the classification results.
The remainder of the paper is organized as follows: section 2 discusses the related work in the literature, section 3 provides a detailed description of the datasets, key characteristics, the features removed, and data ratio, section 4 outlines the problem formulation, the Kolmogorov-Arnold representation theory and KAN, as well as the contrastive training and fine-tuning stages, section 5 first discusses the experimental results, followed by an extensive ablation study investigating KAN’s hyper-parameters, including spline order, grid size and width of the network, visualization of the activation functions learned by the pre-trained feature extractor, and a comparative analysis with other methods are presented as well. Finally section 6 concludes the paper and summarizes the key findings.
@@ 2 Related Works
In [14], different ensemble learning approaches such as XGBoost, random forest, support vector machine (SVM) and stacked DL models, are evaluated for cyber-attack detection. This study highlights the effectiveness of the tree-based models, such as XGBoost in comparison to the DL models. The results indicate that stacked DL model exhibit greater sensitivity to outliers compared to ensemble methods. [15] also utilizes a stacked DL model for detecting cyber-attacks in SCADA systems by combining the outputs of five individual DL models. This approach achieves higher performance compared to individual models. Moreover, this method uses random forest to determine feature importance.
@@
In [16], a CNN-LSTM hybrid network is proposed to improve the temporal dependency in the IIoT networks. However, the authors, acknowledge that the challenges of imbalance data and privacy remain unsolved, and they proposed to adopt federated learning as the future research. A probabilistic hybrid ensemble classifier (PHEC) based on federated learning for IoT networks intrusion detection, is proposed in [17]. While this framework achieves high detection rates, it appears that for the multi-class Gas Pipeline dataset, only a subset of classes was considered, not all classes. This limits fair comparison to the other state-of-the-art methods. Furthermore, the use of order reduction operations in this work, might introduce additional computational complexity. In [18], authors introduce sustainable ensemble learning framework for intrusion detection of industrial control systems (ICS), utilizing sensitivity to different attack types and historical data reuse. In this work, models receive different dynamic importance scores, based on their effectiveness in detecting different types of attacks.
A bidirectional simple recurrent unit (BiSRU) model is introduced in [19] to address the limitations of traditional recurrent neural networks (RNNs), particularly the vanishing gradient problem, using skip connections and bidirectional structure. This BiSRU model improves training efficiency and also enhances the model’s ability to capture forward and backward temporal dependencies. Although this model demonstrate effectiveness and efficiency, its performance is highly sensitive to data balance. [20] proposes an ensemble-based detection model combining deep belief networks (DBNs) with SVMs to address cyber-security challenges, by analyzing network traffic and payload features. One advantage of this method is the one-class training of the SVM model which can be performed with only normal data. However, DBN still requires a large amount of labeled data. Additionally the scalability and training time of SVM for big datasets poses a significant challenge for adoption of this framework in real-world application. The work in [21], focuses on reducing the false positive alarm, by utilizing a graph-guided masked autoencoder (GGMAE) for anomaly detection in the process industry. This approach utilizes a topology graph to simulate process models and applies the Kullback-Leibler divergence as a loss function for input-output consistency.
@@
In [22], contrastive learning with multi-augmentation strategy is employed to encode dynamic graph embeddings. And these embeddings are subsequently passed through transformer-based module to capture evolving patterns. Then a forecasting model predicts long-term trends for anomaly detection. [23] utilizes a feature encoder that converts input sequences into hidden representations, and a neural transformation module that generates multiple latent views. The training is performed using two contrastive losses: a contextual neural contrastive loss that aligns contextual information, and a discriminative contrastive loss that ensures diversity while preserving semantic consistency. In [24], the authors propose an unsupervised hierarchical strategy-based global-local graph contrastive learning (HS-GLGCL) framework to address the challenge of limited labeled data. This approach first constructs a process topology graph based on prior knowledge, then enhances it through graph augmentation and applies graph contrastive learning to extract meaningful representations from the unlabeled data. Finally, anomalies are identified based on reconstruction error thresholds and Kullback-Leibler divergence. [25] introduces an encoder for reinterpretation learning, a decoder for latent reconstruction, a memory network for pattern recognition and a denoiser input correction, which is pre-trained by self-supervised learning. While the approach is comprehensive, the complexity of the framework can hinder real-time intrusion detection.
@@
[26] proposes time-series anomaly detection framework based on learnable data augmentation and contrastive learning in a self-supervised setting. The main innovation of this framework is the learnable data augmentation module that creates challenging negative samples to enhance training. [27] combines contrastive learning with graph neural networks, to learn both similarities in the latent space and dependency structures between different instance pairs in a self-supervised manner.
@@
[28] proposes an improved deviation network with feature selection, to address the need for extensive data for anomaly detection. However, relying on feature selection poses a risk of missing critical attributes, that could potentially lead to undetected intrusions. To address both challenges of limited labeled data and privacy concerns in IIoT applications, [29] utilizes semi-supervised learning integrated with federated learning. Their architecture utilizes a autoencoder with a fully connected layer for intrusions detection. This allows the use of unlabeled data while preserving data privacy. The research in [28] introduces a semi-supervised approach, incorporating information gain (IG) and PCA for feature selection, alongside a Gaussian-based deviation loss function. But feature selection in intrusion detection system, can be cumbersome for intrusion detection, as noted before.
@@
Apart form the learning-based methods, probabilistic and mathematical approaches have also been examined in the literature. One instance is the work proposed in [30], where a probabilistically timed dynamic model for simulating physical security attacks on critical infrastructures is presented. It models an attacker’s behavior in progressing through vulnerabilities. This method calculates mission success probability based on cumulative attack time relative to a given mission duration and uses Monte Carlo sampling to account for uncertainty, and it further produces dynamic attack trees and visual flowcharts, to better represent the decision-making process. [31] employs colored Petri nets and mixed-strategy Nash equilibrium (CPNE) to predict multi-step attack paths for oil and gas intelligent pipelines. This method models asset importance and dependency relationships, and uses this model to simulate attack-defense interactions, then calculates node failure probabilities, extracting multi-step attack paths.
@@ 3 Datasets
In this section, we discuss different industrial control system (ICS) datasets, used in this research. There are plenty of datasets publicly available regarding IoT in the literature, such as SWaT [32], WADI [33]), Gas Pipeline [34], [35], UNSW-NB15 [36] and BoT-IoT [37]. There are two major type of datasets available: simulated datasets such as UNSW-NB15, and datasets collected from physical testbeds, such as SWaT, WADI, Gas Pipeline and BoT-IoT. We validated the results of our method using UNSW-NB15, BoT-IoT and gas pipeline datasets. These datasets are mainly composed of network packets. In the following, we discuss each of the datasets in detail.

### v2

@@ Abstract
In the era of the Fourth Industrial Revolution, cybersecurity and intrusion detection systems are vital for the secure and reliable operation of IoT and IIoT environments. A key challenge in this domain is the scarcity of labeled cyberattack data, as most industrial systems operate under normal conditions. This data imbalance, combined with the high cost of annotation, hinders the effective training of machine learning models. Moreover, the rapid detection of attacks is essential, especially in critical infrastructure, to prevent large-scale disruptions. To address these challenges, we propose a real-time intrusion detection system based on a semi-supervised contrastive learning framework using the Kolmogorov-Arnold Network (KAN). Our method leverages abundant unlabeled data to effectively distinguish between normal and attack behaviors. We validate our approach on three benchmark datasets, UNSW-NB15, BoT-IoT, and Gas Pipeline, using only 2.20%, 1.28%, and 8% of labeled samples, respectively, to simulate real-world conditions. Experimental results show that our method outperforms existing contrastive learning-based approaches. We further compare KAN with a traditional multilayer perceptron (MLP), demonstrating KAN’s superior performance in both detection accuracy and robustness under limited supervision. KAN’s ability to model complex relationships, along with its learnable activation functions, is also explored and visualized, offering interpretability and the potential for rule extraction. The method supports multi-class classification and proves effective in safety, critical environments where reliability is paramount.
@@ 1 Introduction
Today, with the rapid development of networking, the Internet of Things (IoT) and the Industrial Internet of Things (IIoT), including technologies such as smart home applications, medical devices, automated vehicles, and sensor networks [1], are being increasingly adopted. As a result, cybersecurity has become one of the most critical aspects of both cyberspace and network infrastructures. The need for secure digital environments is evident in the era of the Fourth Industrial Revolution, where internet connectivity is integrated into many aspects of modern life [2], [3], [4].
@@
In IoT and IIoT scenarios, intrusion detection systems are designed to identify suspicious patterns in sensor, actuator, and network packet data. With the advent of machine learning (ML) and deep learning (DL), the field of intrusion detection has shifted from classical rule-based methods toward intelligent data-driven approaches. ML techniques require fewer computational resources and less data compared to DL methods. However, they often lack the generalizability needed to handle unseen data. In contrast, DL methods offer superior performance but require large volumes of data and significantly more computational resources [2].
To address these challenges, researchers have proposed various solutions. Unsupervised methods aim to reduce the cost and time associated with data annotation by leveraging unlabeled data [5]. However, these approaches usually fall short in achieving the desired performance. To preserve data privacy, federated learning techniques have been introduced to train models without the need to centralize data [2]. Despite these advances, many existing methods still lack the interpretability required by industry for safe and reliable operation [6].
@@
To address challenges such as data augmentation, labeled data scarcity, and data imbalance, many algorithms rely on oversampling techniques such as the Synthetic Minority Over-sampling Technique (SMOTE) [7] and its variants. Generative methods using Generative Adversarial Networks (GANs) have also been explored. However, there are several issues that limit the effectiveness of such techniques. First, these methods often experience performance degradation when the imbalance ratio is large, which is a common characteristic of intrusion detection datasets. Second, in many of these datasets, the data distributions of attack and normal samples are quite similar. As a result, oversampling techniques can lead to blurred boundaries between attack and normal classes. In the case of GANs, training is often unstable and difficult to converge. In this work, we discourage the use of these methods for handling data imbalance and augmentation. Instead, we adopt a masking strategy that randomly masks out a certain percentage of the data.
@@
Another commonly used technique in intrusion detection is feature selection [8]. However, a significant limitation of this approach is that if an attack targets a specific feature that has been removed during the selection process, the model may fail to detect it. In this work, we avoid using feature selection to preserve potentially important information. The only exception is the removal of features with no variation in the Gas Pipeline dataset. No other form of feature selection is applied.
@@
One of the major challenges in real-time intrusion detection is preprocessing, which can delay the timely detection of cyber intrusions. Many methods use window-based sampling approaches. The problem with such techniques is that the algorithm must wait for a complete data window before proceeding with operations such as averaging or applying one-dimensional convolutional neural networks (1D-CNNs) for feature extraction and final classification. Furthermore, window-based algorithms often exhibit lower performance [9]. Another commonly used preprocessing method is principal component analysis (PCA), which reduces data dimensionality by transforming the input into a set of orthogonal components that capture maximum variance along the principal components. Our approach, in contrast, processes individual data samples, eliminating the need for windowing. The associated preprocessing consists only of feature encoding and standard scaling for normalization. This preserves the interpretability of features and enables timely intrusion detection.
@@
As discussed earlier, the limited availability of annotated data is one of the biggest challenges in machine learning (ML) and deep learning (DL). We have outlined the advantages and disadvantages of both ML and DL. To bridge the gap between these approaches and to address the issue of limited labeled data, semi-supervised methods have been proposed. In these methods, the network is first pretrained on large amounts of unlabeled data, and then fine-tuned using only a small labeled subset. One example of such semi-supervised learning is the SimCLR framework introduced in [10], which we adopt in this work using the Kolmogorov-Arnold Network (KAN). In our study, a feature extractor based on a single layer of KAN is trained using abundant unlabeled data. These data points are first augmented into two different views, and the feature extractor is trained to produce similar representations for these augmented views, which are essentially different versions of the same data point. Once pretraining is completed, fine-tuning is performed using only a small labeled subset.
@@
To address the interpretability and explainability challenges faced by deep learning models, we incorporate the Kolmogorov-Arnold Network (KAN) into our framework. First proposed in [11], KAN is a variant of a fully connected layer where traditional weights are replaced with learnable activation functions represented as splines. These splines approximate univariate functions capable of capturing complex patterns in the data. As will be discussed in later sections, KAN outperforms multilayer perceptron (MLP) networks with the same number of parameters. KAN decomposes complex multivariate inputs into multiple univariate functions. Originally, it was used for solving ordinary and partial differential equations [12], but recent studies have extended KAN’s applications to time series analysis [13] and classification tasks.
The contributions of this study are as follows:
@@
  • An efficient real-time intrusion detection system based on semi-supervised contrastive learning, which leverages large amounts of unlabeled data.
@@
  • Utilizing only a small fraction of labeled data while achieving better performance than other state-of-the-art methods.
@@
  • Conducting a thorough ablation study and comparison with traditional MLP networks, exploring different hyperparameters of the KAN model to identify those that most impact classification results.
The remainder of the paper is organized as follows: section 2 reviews related work in the literature; section 3 provides a detailed description of the datasets, including key characteristics, features removed, and data ratios; section 4 outlines the problem formulation, the Kolmogorov-Arnold representation theory and KAN, as well as the contrastive training and fine-tuning stages. section 5 presents the experimental results, followed by an extensive ablation study investigating KAN’s hyperparameters, including spline order, grid size, and network width. Visualization of the activation functions learned by the pre-trained feature extractor and a comparative analysis with other methods are also included. Finally, section 6 concludes the paper and summarizes the key findings.
@@ 2 Related Works
In [14], different ensemble learning approaches, such as XGBoost, random forest, support vector machine (SVM), and stacked deep learning (DL) models, are evaluated for cyber-attack detection. This study highlights the effectiveness of tree-based models, such as XGBoost, in comparison to DL models. The results indicate that stacked DL models exhibit greater sensitivity to outliers compared to ensemble methods. In [15], a stacked DL model is also utilized for detecting cyber-attacks in SCADA systems by combining the outputs of five individual DL models. This approach achieves higher performance compared to individual models. Moreover, this method employs random forest to determine feature importance.
@@
In [16], a CNN-LSTM hybrid network is proposed to improve temporal dependency modeling in IIoT networks. However, the authors acknowledge that challenges related to data imbalance and privacy remain unresolved, and they suggest adopting federated learning as a direction for future research. A probabilistic hybrid ensemble classifier (PHEC) based on federated learning for IoT network intrusion detection is proposed in [17]. While this framework achieves high detection rates, it appears that, for the multi-class Gas Pipeline dataset, only a subset of classes was considered rather than all classes, limiting fair comparison with other state-of-the-art methods. Furthermore, the use of order reduction operations in this work may introduce additional computational complexity. In [18], the authors introduce a sustainable ensemble learning framework for intrusion detection in industrial control systems (ICS), which utilizes sensitivity to different attack types and historical data reuse. In this framework, models are assigned dynamic importance scores based on their effectiveness in detecting different attack types.
A bidirectional simple recurrent unit (BiSRU) model is introduced in [19] to address the limitations of traditional recurrent neural networks (RNNs), particularly the vanishing gradient problem, by using skip connections and a bidirectional structure. This BiSRU model improves training efficiency and enhances the model’s ability to capture both forward and backward temporal dependencies. Although this model demonstrates effectiveness and efficiency, its performance is highly sensitive to data balance.
[20] proposes an ensemble-based detection model combining deep belief networks (DBNs) with support vector machines (SVMs) to address cyber-security challenges by analyzing network traffic and payload features. One advantage of this method is the one-class training of the SVM model, which can be performed using only normal data. However, DBNs still require a large amount of labeled data. Additionally, the scalability and training time of SVM for large datasets pose significant challenges for the adoption of this framework in real-world applications.
The work in [21] focuses on reducing false positive alarms by utilizing a graph-guided masked autoencoder (GGMAE) for anomaly detection in the process industry. This approach employs a topology graph to simulate process models and applies the Kullback-Leibler divergence as a loss function to ensure input-output consistency.
@@
In [22], contrastive learning with a multi-augmentation strategy is employed to encode dynamic graph embeddings. These embeddings are subsequently passed through a transformer-based module to capture evolving patterns. Then, a forecasting model predicts long-term trends for anomaly detection. [23] utilizes a feature encoder that converts input sequences into hidden representations, and a neural transformation module that generates multiple latent views. Training is performed using two contrastive losses: a contextual neural contrastive loss that aligns contextual information, and a discriminative contrastive loss that ensures diversity while preserving semantic consistency.
In [24], the authors propose an unsupervised hierarchical strategy-based global-local graph contrastive learning (HS-GLGCL) framework to address the challenge of limited labeled data. This approach first constructs a process topology graph based on prior knowledge, then enhances it through graph augmentation and applies graph contrastive learning to extract meaningful representations from the unlabeled data. Finally, anomalies are identified based on reconstruction error thresholds and Kullback-Leibler divergence. [25] introduces an encoder for reinterpretation learning, a decoder for latent reconstruction, a memory network for pattern recognition, and a denoiser for input correction, which is pre-trained by self-supervised learning. While the approach is comprehensive, the complexity of the framework can hinder real-time intrusion detection.
@@
[26] proposes a time-series anomaly detection framework based on learnable data augmentation and contrastive learning in a self-supervised setting. The main innovation of this framework is the learnable data augmentation module that creates challenging negative samples to enhance training. [27] combines contrastive learning with graph neural networks to learn both similarities in the latent space and dependency structures between different instance pairs in a self-supervised manner.
@@
[28] proposes an improved deviation network with feature selection to address the need for extensive data in anomaly detection. However, relying on feature selection poses a risk of missing critical attributes, which could potentially lead to undetected intrusions. To address both the challenges of limited labeled data and privacy concerns in IIoT applications, [29] utilizes semi-supervised learning integrated with federated learning. Their architecture employs an autoencoder with a fully connected layer for intrusion detection, allowing the use of unlabeled data while preserving data privacy. The research in [28] introduces a semi-supervised approach incorporating information gain (IG) and principal component analysis (PCA) for feature selection, alongside a Gaussian-based deviation loss function. However, as noted before, feature selection in intrusion detection systems can be cumbersome.
@@
Apart from learning-based methods, probabilistic and mathematical approaches have also been examined in the literature. One example is the work proposed in [30], which presents a probabilistically timed dynamic model for simulating physical security attacks on critical infrastructures. This model captures an attacker’s behavior as they progress through vulnerabilities. It calculates mission success probability based on cumulative attack time relative to a given mission duration and employs Monte Carlo sampling to handle uncertainty. Additionally, it generates dynamic attack trees and visual flowcharts to better represent the decision-making process.
In [31], colored Petri nets and mixed-strategy Nash equilibrium (CPNE) are used to predict multi-step attack paths for oil and gas intelligent pipelines. This method models asset importance and dependency relationships, simulates attack-defense interactions, and calculates node failure probabilities, from which multi-step attack paths are extracted.
@@ 3 Datasets
In this section, we discuss the IoT datasets used in this research. Numerous publicly available IoT datasets exist in the literature, such as SWaT [32], WADI [33], Gas Pipeline [34, 35], UNSW-NB15 [36], and BoT-IoT [37].
There are two major types of datasets: simulated datasets, such as UNSW-NB15, and datasets collected from physical testbeds, such as SWaT, WADI, Gas Pipeline, and BoT-IoT. We validated our method using the UNSW-NB15, BoT-IoT, and Gas Pipeline datasets. These datasets primarily consist of network packet data. In the following, we provide a detailed discussion of each dataset.